13. Order: Sustacal 240 mI via PEG, infuse in 60 minutes At what rate will you set the infusion pump? 14. A patient is receiving an IV of 1 NS at 125 mL/h. There are 650 mL in the IV bag at 1450 h. What time will the infusion be completed?

media%2F739%2F739ed409-3ff4-45ff-bb02-d2
Answer #1

13. Sustacal =240ml(Volume)

Time= 60 minutes = 1 hour

Since the volume is given in mL and the time is given in hours, the flow rate can be calculated in one step using the formula.

Flow rate (mL/Hr )= Volume(mL)/Time(Hr)

= 240 /1=240 ml /hr

Flow rate (mL/min)= Volume(mL)/Time(minutes)

=240/60= 4mL/min

-------------------------------------------------------------------------------------------------------------------------------------------------------------------------------

14. IV fluid is infused at rate of 125 mL/Hr to a patient.

At 1450H 650 ml fluid is remaining in the bag

So at 1450H= 650 mL

at 1550H= 525mL(650-125)

at 1650H=400mL(525-125)

at 1750H=275mL(400-125)

at 1850H=150mL(275-125)

at 1950H=25mL(150-125)

The fluid infused mL/min= 125/60= 2.08mL/min = 2 mL /min( rounded as drops cannot be in decimals)

25 mL is infused in 12 minutes ( 12mins* 2ml= 24 ml)

So at 2002H infusion may get ended and approximated to 2000H

--------------------------------------------------------------------------------------------------------------------------------------------------------------------------------

16. Ceftazidime 50mg/Kg is given to a patient weighing 40 kg.

a) Weight (kg) x Dosage Ordered (per kg) = Y (Required Dosage)

weight = 40 kg, Dosage ordered= 50mg/kg

The patient is receiving 2g Q8H .( 40 kg * 50mg = 2000 mg= 2g)

b) Time = 30 minutes, Volume = 100ml , drop factor = 10 gtts/ml

Volume (mL) /Time (min) x Drop Factor (gtts/mL) = Y (Flow Rate in gtts/min)

100/30*10=33.33

Flow rate in drops / minute = 33 drops/min ( rounded as drops cannot be in decimals)